# A Computationally Efficient Maximum A Posteriori Sequence Estimation via Stein Variational Inference

### v3

@@ V-B Pose Estimation under Unknown Data Association
To evaluate the state estimation methods, we generated $10$ independent trajectories of $630$ time steps each. For each estimator, we computed the RMSE of the state estimates, averaged over all Monte Carlo runs. The results in Fig. 4, Fig. 5, and Table II demonstrated that Stein-MAP-Seq significantly outperformed several state estimation methods, indicating that Stein-MAP-Seq is robust to multimodal posterior distributions.
Gaussian-based methods, such as EKF, EKS, iEKF-3, and iEKS-3, were not robust to multi-peak posterior distributions. To assess the effect of initialization in iterative methods, we also implemented EKS-GT and iEKS-3-GT, both initialized at the ground truth. Monte Carlo results for EKS-GT were generally close to the ground truth (see Fig. 4, 1st quadrant). In contrast, iEKS-3-GT produced worse estimates than EKS-GT, with performance degrading further as the number of iterations increased. This degradation occurs because each re-linearization step perturbs the linearization point due to process and measurement noise, potentially causing the estimate to drift away from the true state in nonlinear and multimodal settings.
Among PF-based methods, PF-MAP with $500$ particles achieved the best performance (Table II and Fig. 5). PF-MAP outperformed PF, an MMSE estimator, across all tested particle counts. In contrast, PF-MAP-Seq exhibited mode switching, resulting in a zig-zag trajectory and the poorest estimation performance (see Fig. 4, 3rd quadrant). We also observed that, even as the number of particles increased, the estimation performance of PF-based methods tended to degrade. This indicates that more particles do not necessarily provide a better representation of the posterior. Even with importance sampling, randomly drawn particles often populate intermediate-probability regions (neither high nor low), which reduces their concentration around the true mode and consequently degrades estimation accuracy.
Stein-series methods outperformed both Gaussian-based and PF-based methods. In particular, Stein-MAP-Seq with $50$ particles achieved the best overall performance (Table II). For SPF, SPF-MAP, and Stein-MAP-Seq, estimation performance improved with increasing particle count up to $50$.
TABLE II: Robustness Evaluation for Scenario (B), Averaged Pose RMSE
EKF EKS EKS-GT $N_{s}$ PF PF-MAP PF-MAP-Seq
5e2 $0.1818$ $\mathbf{0.1745}$ $0.2240$
$2.1028$ $2.4934$ $\mathbf{0.0832}$ 1e3 $0.1845$ $0.1787$ $0.2313$
2e3 $0.1872$ $0.1812$ $0.2442$
iEKF-3 iEKS-3 iEKS-3-GT $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
2e1 $0.0514$ $0.0522$ $0.0510$
$2.1070$ >100>100 >100>100 4e1 $0.0382$ $0.0375$ $0.0394$
5e1 $0.0299$ $0.0298$ $\mathbf{0.0293}$
@@ V-C Range-only Localization under Temporary Unobservability
In the last example, we consider a range-only localization task [43, 44]. In our experiment, a target equipped with a DWM1000 Ultra-wideband (UWB) transceiver (tag) traversed along an arbitrary trajectory in an 88\,m $\times~8\,$m indoor environment with 3 UWB anchors. Here, UWB provides range measurements at 10 Hz and can also communicate their IDs, eliminating the need for data association considerations. There were no obstacles between the anchors and the tag to mitigate the biased range measurements. For a precise reference trajectory comparison, we employed the OptiTrack optical motion capture system with 12 infrared cameras to achieve high precision localization with an accuracy of $10^{-4}\;$m, and a sampling rate of $120$ Hz.
We localize a target that is not equipped with any proprioceptive sensor, such as an IMU or encoders. Consequently, we adopt a zero-velocity motion model, given by
@@
where the process noise is modeled as $\mathbf{v}_{t-1}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{t-1})$, with $\mathbf{Q}_{t-1}=\mathrm{diag}\bigl{[}(\alpha_{x}\,\Delta t)^{2},\;(\alpha_{y}\,\Delta t)^{2}\bigr{]}$. The state vector $\mathbf{x}_{t}=\bigl{[}x_{t},\;y_{t}\bigr{]}\in{\mathbb{R}}^{2}\,(m)$ represents the 2D position at time $t$. We use line-of-sight (LoS) range measurements of the form:
@@
$p(\mathbf{z}_{t}|\mathbf{x}_{t})=\prod_{\forall l}p(z_{t,l}|\mathbf{x}_{t}),$ (28)
where the product arises from the Assumption II.2. In this experiment, we set $\alpha_{x}=\alpha_{y}=1.0$, $\Delta t=0.1$, and $\sigma_{z}=0.5$ to model measurement noise and compensate for minor range bias.
In this scenario, some anchors may be blocked by dynamic or static obstacles, leading to a multimodal distribution due to the underdetermined nature of distance measurements. In the experiment, we dropped measurements from anchor #1 at specific time instances to simulate real-world (underdetermined) scenarios for robustness analysis against a symmetric multimodality, see Fig 6.
(a) Blocking #1
(b) Blocking #2
(c) Blocking #3
(d) Blocking #4
(e) Blocking #5
(f) Blocking #6
Figure 7: Estimation results at $10$ Hz during each blocking period. Stein-MAP-Seq achieved robust estimation across all multimodal periods.
@@
First, we observed that Gaussian-based methods, such as EKF and iEKF-2, often converged to an incorrect mode during most blocking periods. This outcome was examined by analyzing the contour plot of the posterior distribution shown in Fig. 8. To assess the effect of initialization in smoothing or iterative methods, we also implemented EKS-GT and iEKS-3-GT, both initialized with high-precision position data. EKS-GT generally produced estimates close to the high-precision position data due to its locally optimal convergence. In contrast, iEKS-3-GT yielded worse estimates than EKS-GT, with performance degrading further as the number of iterations increased, similar to the position estimation results in Subsection V-B.
In contrast, the PF method can manage non-Gaussian distributions using weighted particles. Nonetheless, it had limitations in fully addressing symmetric multimodality. This is because PF employs the transition model as a proposal distribution for importance sampling, resulting in particle weights that reflect the inherently symmetric, multimodal nature of the likelihood (see Fig. 8). While PF-MAP achieved slightly better performance than PF in handling multimodality—yielding a lower RMSE at certain points—it still could not resolve the symmetric multimodality, resulting in overall performance only marginally better than PF. For PF-MAP-Seq, the symmetric multimodality caused mode switching, which produced worse estimation results than both PF and PF-MAP.
In the case of SPF and SPF-MAP, despite using only 200200400400 particles, they outperformed both PF and PF-MAP with 1000100040004000 particles in terms of accuracy. However, SPF and SPF-MAP also had limitations in handling symmetric multimodality. As shown in Fig. 8, the particles were positioned according to the gradient flow, allowing a relatively small number of particles to represent the multimodal distribution; however, the symmetric nature of the distribution still led to low estimation accuracy. Table III and Fig. 7 demonstrated that our proposed Stein-MAP-Seq consistently maintains robustness against symmetric multimodal posteriors compared with the other methods.
TABLE III: Robustness Evaluation for Scenario (C), RMSE (m)
EKF EKS EKS-GT $N_{s}$ PF PF-MAP PF-MAP-Seq
@@
iEKF-2 iEKS-3 iEKS-3-GT $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
2e2 $0.1933$ $0.1908$ $0.1316$
$0.1981$ $0.1921$ $0.1243$ 3e2 $0.1932$ $0.1955$ $\mathbf{0.1273}$
4e2 $0.1931$ $0.1934$ $0.1307$
@@
Figure 8: Contour plots of the Posterior for each method during Blocking #1 and #2 periods.

### v4

@@ V-B Pose Estimation under Unknown Data Association
To evaluate the state estimation methods, we generated $10$ independent trajectories of $630$ time steps each. For each estimator, we computed the RMSE of the estimated states with respect to the noise-free reference trajectory and averaged the results over all Monte Carlo runs. The results in Fig. 5, Fig. 6, and Table III demonstrated that Stein-MAP-Seq consistently outperformed Gaussian-assumed methods, PF-based methods, and batch MAP estimation. This performance highlighted its strong robustness to highly multimodal and heavy-tailed posterior distributions arising from unknown data association with outliers.
TABLE III: Robustness Evaluation for Scenario (B), Averaged Pose RMSE
Gaussian-assumed Estimation Finite-state Estimation
EKF EKS EKS-GT GTSAM $N_{s}$ PF PF-MAP PF-MAP-Seq
5e2 $0.5609$ $0.5638$ $0.5825$
$0.5903$ $0.5442$ $\mathbf{0.0179}$ $0.1520$ 1e3 $\mathbf{0.5337}$ $0.5370$ $0.5512$
2e3 $0.6223$ $0.6239$ $0.6457$
iEKF-3 iEKS-3 iEKS-3-GT GTSAM-Stein $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
1e1 $0.1146$ $0.1142$ $0.1121$
$0.6259$ $0.6392$ $0.0769$ $0.1022$ 2e1 $0.0996$ $0.0993$ $\mathbf{0.0971}$
4e1 $0.0996$ $0.0989$ $0.0983$
As illustrated in Fig. 5, (b) and Fig. 6, Gaussian-based estimators exhibited limited robustness in the presence of multi-peak posterior distributions. In particular, these methods tended to collapse onto a single mode, resulting in biased trajectory estimates. To isolate the effect of initialization in iterative smoothers, we additionally evaluated EKS-GT and iEKS-3-GT, both initialized at the noise-free ground-truth trajectory, as shown in Table III. While EKS-GT achieved accurate estimates, iEKS-3-GT produced slightly degraded results as the number of iterations increased. This behavior arose because repeated re-linearization around incorrect linearization points caused the estimate to drift toward spurious local modes.
As shown in Fig. 5, (b) and Table III, GTSAM improved trajectory smoothness by jointly optimizing the entire trajectory. In this simulation, since GTSAM does not support association-marginalized likelihoods, we adopted a nearest-landmark association strategy based on the observed measurement $\mathbf{z}_{t}$, rather than employing robust loss functions. Under this setting, GTSAM exhibited noticeable bias because high process uncertainty weakened the constraints imposed by the motion model, causing the estimator to rely more heavily on measurements. Consequently, when data associations were incorrect or measurements contained large outliers, these effects dominated the nonlinear least-squares objective. As a result, the optimizer often converged to a suboptimal local minimum that explained the measurements. These results indicated that the robustness of batch MAP estimation was limited in long-horizon localization problems characterized by high process uncertainty and outlier-contaminated observations.
As shown in Fig. 5, (c), PF-based estimators improved robustness compared to Gaussian-assumed methods by explicitly representing multiple hypotheses. With a sufficiently large number of particles, PF-based estimators were able to recover the trajectory and mitigate severe bias. However, despite this improvement, PF-based methods still exhibited noticeable trajectory distortion due to heavy-tailed outliers. This effect progressively reduced particle diversity over long horizons and led to premature mode collapse. Consequently, even with large particle sets, as reported in Table III, PF-based estimators struggled to maintain a coherent global trajectory, highlighting fundamental limitations in robustness.
As shown in Fig. 5, (d), Stein-based estimators consistently recovered the reference trajectory with high accuracy. Unlike PF-based methods, Stein-based approaches maintained a compact yet expressive particle set, effectively preserving multiple modes. This property enabled Stein-based estimators to avoid particle degeneracy and the associated trajectory distortions. As shown in Table III, accurate trajectory reconstruction was achieved with a substantially smaller number of particles, demonstrating both robustness and computational efficiency. These results highlighted the advantage of Stein-MAP-Seq in maintaining global trajectory consistency under persistent multimodality, making it particularly well-suited for long-horizon state estimation problems characterized by ambiguous measurements and heavy-tailed outliers.
Figure 7: The averaged computation time per step is reported over $10$ Monte Carlo simulations, each consisting of $630$ time steps. Gaussian filtering and smoothing methods exhibit sub-millisecond runtimes due to their closed-form updates. GTSAM incurs a higher per-step computational cost than filtering–smoothing methods. PF-based MAP methods scale with the number of particles, resulting in rapidly increasing computation times due to forward recursion over large particle sets. Stein-based methods exhibit moderate computational costs by relying on a relatively small number of particles.
Alongside estimation accuracy, Fig. 7 compared the per-step computational cost of the considered methods, revealing a clear accuracy–efficiency trade-off. Consistent with previous scenario, in the 3D case, Gaussian-assumed methods achieved the lowest computational cost due to closed-form updates and local temporal processing, but exhibited limited robustness under multimodal uncertainty. GTSAM incurred higher computational cost than filtering–smoothing methods, which improved robustness but remained sensitive to long-horizon multimodality. PF-based MAP estimators exhibited a rapid increase in computation time as the particle count increased and failed to maintain global trajectory coherence despite the higher cost. In contrast, Stein-MAP-Seq achieved strong robustness with moderate computational cost using significantly fewer particles, offering a favorable accuracy–efficiency trade-off in long-horizon, multimodal settings.
@@ V-C Range-only Localization under Temporary Unobservability
In the third example, we consider a range-only localization task [47, 48]. In our experiment, a target equipped with a DWM1000 Ultra-wideband (UWB) transceiver (tag) traversed along an arbitrary trajectory in an 88\,m $\times~8\,$m indoor environment with 3 UWB anchors. Here, UWB provides range measurements at 10 Hz and can also communicate their IDs, eliminating the need for data association considerations. There were no obstacles between the anchors and the tag to mitigate the biased range measurements. For a precise reference trajectory comparison, we employed the OptiTrack optical motion capture system with 12 infrared cameras to achieve high precision localization with an accuracy of $10^{-4}\;$m, and a sampling rate of $120$ Hz.
Figure 9: Estimation results at $10$ Hz across 66 blocking periods. EKF, PF, SPF, and GTSAM exhibit trajectory fragmentation at certain blocking intervals due to likelihood-driven updates under symmetric multimodal ambiguity. In contrast, Stein-MAP-Seq maintains stable MAP sequence estimates across all blocking periods.
We localize a target that is not equipped with any proprioceptive sensor, such as an IMU or wheel encoders, making the localization problem particularly challenging due to the absence of motion priors. Consequently, we adopt a zero-velocity motion model, given by
@@
where the process noise is modeled as $\mathbf{v}_{t-1}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{t-1})$, with $\mathbf{Q}_{t-1}=\mathrm{diag}\bigl[(\alpha_{x}\,\Delta t)^{2},\;(\alpha_{y}\,\Delta t)^{2}\bigr]$. The state vector $\mathbf{x}_{t}=\bigl[x_{t},\;y_{t}\bigr]^{\top}\in{\mathbb{R}}^{2}\,(m)$ represents the 2D position at time $t$. We use line-of-sight (LoS) range measurements of the form:
@@
$p(\mathbf{z}_{t}|\mathbf{x}_{t})=\prod_{\forall l}\nolimits p(z_{t,l}|\mathbf{x}_{t}),$ (28)
In this experiment, we set $\alpha_{x}=\alpha_{y}=1.0$, $\Delta t=0.1$, and $\sigma_{z}=0.5$ to model measurement noise and compensate for minor range bias.
In this scenario, some anchors may be blocked by dynamic or static obstacles, leading to a multimodal distribution due to the underdetermined nature of distance measurements. In the experiment, we dropped measurements from anchor #1 at specific time instances to simulate real-world scenarios for robustness analysis against a symmetric multimodality, see Fig 8.
@@
This scenario evaluated range-only localization under temporary unobservability, where measurements were intermittently unavailable, resulting in periods of weak observability and unknown motion. During these intervals, similar to previous scenarios, Gaussian-assumed methods quickly diverged due to linearization errors and their inability to represent multiple hypotheses.
In this scenario, GTSAM exhibited worse performance than filtering–smoothing methods. This behavior arose from the absence of effective motion constraints, combined with a lack of informative measurement factors during periods of temporary unobservability. As a result, GTSAM often converged to globally incorrect trajectories and failed to recover even after observability was restored.
PF-based methods exhibited different failure modes. PF methods relied on prior-based proposals with likelihood-weighted importance sampling, as shown in Fig. 10, which caused particles to prematurely collapse onto locally dominant but globally incorrect modes during unobservable intervals. Once particles were depleted, recovery remained difficult even after measurements became available again. Consequently, PF-MAP-Seq produced worse estimation results than both PF and PF-MAP due to frequent mode switching.
In contrast, Stein-MAP-Seq maintained multiple MAP-consistent hypotheses during unobservable periods. By conditioning on the previous MAP estimates, the proposed method remained robust to temporary loss of observability and consistently recovered the correct trajectory once measurements were reintroduced. Table IV and Fig. 9 demonstrated that Stein-MAP-Seq consistently maintained robustness against symmetric multimodal posteriors compared with the other methods.
Regarding computational cost, as shown in Table V, GTSAM incurred increased computation time due to a larger number of optimization iterations caused by the absence of informative motion priors and observations during the blocking period. In contrast, Stein-MAP-Seq achieved a moderate and predictable computational cost, consistent with its performance in the previous scenario, and outperformed other MAP-Seq estimators in terms of efficiency.
TABLE IV: Robustness Evaluation for Scenario (C), RMSE (m)
Gaussian-assumed Estimation Finite-state Estimation
EKF EKS EKS-GT GTSAM $N_{s}$ PF PF-MAP PF-MAP-Seq
@@
iEKF-2 iEKS-3 iEKS-3-GT GTSAM-Stein $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
2e1 $0.1921$ $0.1969$ $0.1722$
$0.1981$ $0.1921$ $0.1243$ $0.1228$ 3e1 $0.1925$ $0.1902$ $0.1442$
4e1 $0.1949$ $0.1858$ $\mathbf{0.1369}$
TABLE V: Computation Time Evaluation for Scenario (C), (msec)
Gaussian-assumed Estimation Finite-state Estimation
EKF EKS GTSAM $N_{s}$ PF PF-MAP PF-MAP-Seq
1e3 $2.463$ $53.23$ $58.94$
$0.1458$ $0.1531$ $36.69$ 2e3 $2.689$ $148.12$ $162.01$
4e3 $3.234$ $460.36$ $505.32$
iEKF-2 iEKS-3 GTSAM-Stein $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
2e1 $9.116$ $9.736$ $10.82$
$0.1905$ $0.4505$ $13.27$ 3e1 $11.60$ $12.56$ $13.60$
4e1 $14.04$ $15.28$ $16.44$
@@
Figure 10: Posterior behaviors of EKF, PF, SPF and Stein-MAP-Seq during Blocking #1 and #2 periods. EKF collapses the posterior to a single incorrect mode due to local linearization, while PF relies on likelihood-based importance sampling, leading to premature mode commitment. SPF improves multimodal posterior representation but cannot handle symmetric uncertainty. In Stein-MAP-Seq, the displayed particles do not represent posterior samples; instead, they visualize score-informed particle locations arising from the forward recursion toward a posterior defined conditionally on the previous MAP estimates $\mathbf{x}_{1:t-1}^{\star}$, thereby enabling accurate MAP sequence estimation.
V-D High-Dimensional Manipulator with Kinematic Redundancy
In the last example, we consider a high-dimensional nonlinear system using a $7$-DoF robotic manipulator. The system state consists of joint positions and velocities, yielding a $14$-dimensional state vector, given by
$\mathbf{x}_{t}=\begin{bmatrix}\mathbf{q}_{t}^{\top}&\dot{\mathbf{q}}_{t}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{14},$ (29)
where $\mathbf{q}_{t},\dot{\mathbf{q}}_{t}\in\mathbb{R}^{7}$ denote the joint angles and joint angle velocities, respectively.
The manipulator dynamics are modeled using a discrete-time second-order system with sampling period $\Delta t=0.02$,
$\mathbf{q}_{t}$ $=\mathbf{q}_{t-1}+\dot{\mathbf{q}}_{t-1}\Delta t,$ (30a)
$\dot{\mathbf{q}}_{t}$ $=\dot{\mathbf{q}}_{t-1}+\mathbf{u}_{t-1}\Delta t+\mathbf{v}_{t-1},$ (30b)
where $\mathbf{u}_{t-1}\in\mathbb{R}^{7}$ is the joint acceleration input and $\mathbf{v}_{t-1}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{t-1})$ denotes Gaussian process noise. The covariance is chosen block-diagonal as $\mathbf{Q}_{t-1}=\mathrm{diag}(q_{p}\Delta t^{2}\mathbf{I}_{7},~q_{v}\mathbf{I}_{7})$ with $q_{p}=1.0$ and $q_{v}=5\times 10^{-3}$. The control input is a smooth joint-wise sinusoid, $u_{t}^{(i)}=0.5\sin(0.2t+\phi_{i}),~~i=1,\dots,7$ where $\phi_{i}=i-1$ introduces fixed phase offsets across joints, resulting in rich excitation and strong nonlinear coupling in the forward kinematics.
At each time step, the system receives an exteroceptive observation of the end-effector position,
$\mathbf{z}_{t}=\mathbf{h}(\mathbf{q}_{t})+\mathbf{r}_{t},$ (31)
where $\mathbf{h}(\cdot)$ denotes the forward kinematics mapping and $\mathbf{r}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ with $\mathbf{R}=(0.1)^{2}\mathbf{I}_{3}$. The forward kinematics are defined by a Denavit–Hartenberg chain, yielding a nonlinear and kinematically redundant mapping from joint space to task space, which naturally induces multimodal posteriors in joint space as shown in Fig. 11.
The initial state is distributed as $\mathbf{x}_{0}\sim\mathcal{N}(\boldsymbol{\mu}_{0},\boldsymbol{\Sigma}_{0})$, where $\boldsymbol{\mu}_{0}$ initializes joint angles uniformly in [0.5,0.5][-0.5,0.5] (rad) with zero velocities. The initial covariance is block-diagonal, $\boldsymbol{\Sigma}_{0}=\mathrm{diag}(\sigma_{q}^{2}\mathbf{I}_{7},~\sigma_{\dot{q}}^{2}\mathbf{I}_{7})$, with $\sigma_{q}=1^{\circ}$ and $\sigma_{\dot{q}}=2^{\circ}$.
Figure 11: Ground-truth end-effector trajectory in task space $\mathbb{R}^{3}$ (top left), together with measurement log-likelihood surfaces over the $(q_{1},q_{2})$ subspace at selected time steps (bottom and top right). Despite smooth motion in task space, the likelihood landscapes in joint space exhibit broad ridge-like structures and multiple local maxima, reflecting strong kinematic redundancy and multimodality.
Figure 12: End-effector position RMSE over 20 Monte Carlo runs for all compared methods. All methods exhibit noticeable run-to-run variability due to kinematic redundancy and long-horizon error accumulation. Despite this variability, the Stein-based estimators achieve lower RMSE than the other methods.
Performance is evaluated using joint-space and end-effector RMSE, averaged over $20$ Monte Carlo runs. In this scenario, the $7$-DoF manipulator induces a highly multimodal posterior in joint space due to kinematic redundancy and end-effector–only observations. As illustrated in Fig. 11, smooth and well-behaved motion in task space gives rise to broad, ridge-like likelihood structures with multiple local maxima in joint space, making state estimation particularly challenging.
Quantitative results in Fig. 12 and Table VI showed that Stein-MAP-Seq consistently achieved the lowest end-effector RMSE among all compared methods. All methods exhibited large variance because multiple joint configurations were consistent with the same end-effector measurement, resulting in strong multimodality. Unlike previous scenarios, GTSAM showed inferior performance under kinematic redundancy.
PF-based estimators required a substantially larger number of particles to maintain trajectory coherence in this $14$-dimensional state space. As shown in Fig. 12, even with an increased particle count, PF-based MAP estimators suffered from degraded accuracy, reflecting particle path degeneracy and unstable mode selection. In contrast, Stein-MAP-Seq achieved superior performance using as few as 10103030 particles, highlighting the advantage of mode-seeking behavior combined with forward recursion in maintaining mode-coherent support for MAP sequence estimation, even in high-dimensional scenarios.
Joint-space RMSE results in Table VII further revealed an important distinction between task-space and configuration-space accuracy. Even when comparable end-effector RMSE was achieved, most estimators exhibited significantly larger joint-space errors, indicating inconsistency in resolving kinematic redundancy. In contrast, Stein-MAP-Seq achieved the lowest joint-space RMSE among MAP-Seq estimators, demonstrating its ability to maintain consistency in both task space and configuration space. Moreover, the trajectory estimates produced by Stein-MAP-Seq provided effective initialization for GTSAM, guiding the optimizer toward correct joint-space modes and thereby improving configuration-space estimation accuracy.
Finally, the computational comparison in Fig. 13 showed that Stein-MAP-Seq achieved a favorable accuracy–efficiency trade-off. Despite operating in a $14$-dimensional state space, a small number of particles was sufficient to represent dominant modes, thereby maintaining coherent MAP trajectory hypotheses. Moreover, even as the state dimension $n_{x}$ increased, particle updates were performed in parallel on computing hardware, maintaining computational cost comparable to that of lower-dimensional scenarios. Consequently, the runtime remained comparable to GTSAM and was substantially lower than that of PF-based MAP estimators. These results demonstrated that Stein-MAP-Seq scaled effectively to high-dimensional, kinematically redundant systems and provided robust MAP-Seq estimation.
TABLE VI: Robustness Evaluation for Scenario (D), End-effector RMSE (m)
Gaussian-assumed Estimation Finite-state Estimation
EKF EKS EKS-GT GTSAM $N_{s}$ PF PF-MAP PF-MAP-Seq
5e2 $0.6115$ $0.6134$ $0.6120$
$0.6078$ $0.6007$ $\mathbf{0.0617}$ $0.9198$ 1e3 $0.5946$ $0.5912$ $0.5789$
2e3 $0.5409$ $0.5393$ $\mathbf{0.5370}$
iEKF-3 iEKS-3 iEKS-3-GT GTSAM-Stein $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
1e1 $0.4712$ $0.4708$ $0.4686$
$0.6087$ $0.6087$ $0.1242$ $0.5647$ 2e1 $0.4652$ $0.4652$ $0.4624$
3e1 $0.3584$ $0.3584$ $\mathbf{0.3560}$
TABLE VII: Robustness Evaluation for Scenario (D), Joint Angles $\mathbf{q}_{t}$ RMSE (rad)
Gaussian-assumed Estimation Finite-state Estimation
EKF EKS EKS-GT GTSAM $N_{s}$ PF PF-MAP PF-MAP-Seq
5e2 $0.8258$ $0.9622$ $0.9581$
$0.8534$ $0.9132$ $\mathbf{0.1334}$ $4.4841$ 1e3 $0.7792$ $0.7942$ $0.8465$
2e3 $\mathbf{0.7329}$ $0.7713$ $0.8668$
iEKF-3 iEKS-3 iEKS-3-GT GTSAM-Stein $N_{s}$ SPF SPF-MAP Stein-MAP-Seq
1e1 $0.1922$ $0.1921$ $0.1918$
$0.8668$ $1.0430$ $0.2320$ $0.9636$ 2e1 $0.1848$ $0.1848$ $0.1843$
3e1 $0.1728$ $0.1728$ $\mathbf{0.1720}$
V-E Discussion on Empirical Performance and Design Insights of Stein-MAP-Seq
This subsection discusses the empirical performance of Stein-MAP-Seq with a focus on the role of kernel choice, particle count, and kernel hyperparameters. Beyond reporting estimation accuracy, our goal is to provide practical insights into why Stein-MAP-Seq performs favorably in multimodal MAP-Seq estimation, clarify its limitations, and discuss practical considerations arising from modeling assumptions in real robotic systems.
V-E1 Effect of Kernel Choice
The kernel function plays a fundamentally different role in MAP-Seq estimation than in MMSE-based estimation. While MMSE-oriented methods aim to approximate the full posterior distribution and are therefore sensitive to kernel-induced distributional shaping, MAP-Seq estimation primarily benefits from coherent particle concentration around dominant modes rather than uniform posterior coverage. To assess the impact of kernel choice, we conducted an ablation study comparing RBF, IMQ, and Matérn ($\nu=2/3$) kernels [49, 50], all of which preserve the same parallelizable pairwise interaction structure while introducing heavy-tailed repulsion (IMQ) and more localized particle interactions with finite correlation length (Matérn), respectively. As reported in Table VIII, performance differences across kernels are relatively small, indicating that kernel shape alone is not a dominant factor in MAP-Seq estimation.
Figure 13: The averaged computation time per time step is reported over $20$ Monte Carlo simulations, each consisting of $250$ time steps. GTSAM exhibits variable computation time due to differences in convergence behavior across runs. Stein-MAP-Seq, despite operating in a high-dimensional state space, achieves comparable computation time by using a small number of particles and fully parallelizable particle updates. In contrast, PF-based MAP sequence estimators incur the highest computational cost as a result of their reliance on large particle counts.
V-E2 Estimation Accuracy and Particle Count
Unlike MC sampling-based MAP-Seq estimators, Stein-MAP-Seq does not exhibit monotonic performance improvement with increasing particle count. Empirically, a relatively small number of particles (typically 105010-50), when combined with an appropriate kernel bandwidth, suffices to achieve accurate MAP trajectory recovery across all tested scenarios. As the particle count increases, the repulsive interaction induced by the kernel term in Eq. (IV.3) becomes more pronounced, potentially spreading particles into lower-density regions of the state space. While such diversity is advantageous for MMSE-based estimation, it can adversely affect MAP-Seq estimation by increasing path ambiguity during the forward recursion in Eq. (21). Consequently, excessive particle counts may reduce trajectory coherence and lead to increased estimation error. This trend is quantitatively supported by Table VIII, which shows that increasing the particle count yields only marginal performance improvements or even degrades estimation accuracy. These results highlight a fundamental distinction between Stein-MAP-Seq and MC sampling-based MAP-Seq estimators, which typically rely on large particle sets.
V-E3 Kernel Bandwidth and Particle Count
Kernel bandwidth controls the trade-off between mode concentration and separation by adjusting particle repulsion. Smaller bandwidths yield more concentrated particle clusters, whereas larger bandwidths induce stronger repulsive interactions and improved mode separation. MAP-Seq estimation is therefore more sensitive to bandwidth choice than MMSE-based estimation, reflecting its mode-seeking nature, where trajectory decoding depends on temporally consistent relative scores rather than marginal posterior coverage. These trends are quantitatively confirmed in Table VIII. For the RBF kernel, increasing the bandwidth from $0.5h$ to $3h$ substantially reduces RMSE even with fewer particles, while smaller bandwidths require larger particle sets to achieve comparable performance. Overall, kernel bandwidth and particle count jointly shape trajectory-level score accumulation, providing practical design guidance for MAP-Seq estimation.
TABLE VIII: Ablation study on kernel type, kernel bandwidth, and particle count. Results are averaged over Monte Carlo runs in Scenario A. Unless otherwise specified, median heuristic bandwidth $h$ and $N_{s}=10$ are used.
Ablation Factor Kernel Setting RMSE
Kernel type RBF $N_{s}=10$ 2.2643
Kernel type IMQ $N_{s}=10$ 2.2512
Kernel type Matérn $N_{s}=10$ 2.2344
Bandwidth scale RBF $0.5h,N_{s}=10$ 2.4887
Bandwidth scale RBF $3h,N_{s}=10$ 2.0462
Bandwidth scale & Particle count RBF $0.5h,N_{s}=40$ 2.2507
Bandwidth scale & Particle count RBF $3h,N_{s}=40$ 2.1970
Particle count RBF $N_{s}=20$ 2.2108
Particle count RBF $N_{s}=40$ 2.1369
Particle count IMQ $N_{s}=20$ 2.1557
Particle count IMQ $N_{s}=40$ 2.1780
Particle count Matérn $N_{s}=20$ 2.1232
Particle count Matérn $N_{s}=40$ 2.0714
V-E4 Practical Considerations Beyond Ideal Smoothness and Independence Assumptions
Stein-MAP-Seq assumes locally smooth system dynamics and conditionally independent observations; however, these assumptions are often violated in real robotic systems due to nonsmooth or discontinuous dynamics and temporally correlated high-rate sensor measurements. Importantly, Stein-MAP-Seq does not rely on global smoothness or strict observation independence. Instead, it operates through local gradient-based particle updates and trajectory-level score accumulation, making it tolerant to moderate deviations from these assumptions, with locally defined or surrogate gradients being sufficient for MAP-Seq estimation. In practice, local nonsmoothness typically affects only short time intervals and does not invalidate the overall MAP trajectory as long as the dominant mode remains temporally consistent, while mild observation correlations mainly change how confident the estimate is, rather than where the mode is. Such effects often manifest as noisier gradients or increased uncertainty and can be mitigated through kernel bandwidth adjustment and particle interaction regularization. As with classical filtering and batch MAP methods, extreme violations may degrade performance, and characterizing behavior under severe nonsmooth dynamics and strongly correlated observations remains an important direction for future work.
V-E5 Complementarity with Batch MAP Optimization
The comparison with batch MAP optimization using GTSAM reveals complementary strengths. While GTSAM jointly optimizes the entire trajectory, its performance is sensitive to initialization under strongly multimodal and nonconvex objectives. In our demonstrations, initializing GTSAM with the Stein-MAP-Seq trajectory guided the optimizer toward the correct mode, resulting in faster convergence and improved estimation accuracy. These results suggest that Stein-MAP-Seq can serve not only as a standalone MAP-Seq estimator, but also as an effective global mode-discovery front-end for multimodal batch MAP optimization.